ELECTRONIC SPECTROSCOPY OF THE JET-COOLED HPBr FREE RADICAL

\underline{BRANDON S. TACKETT} and DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The Ã^{2}A′ - X̃^{2}A′′ electronic transition of the HPBr free radical has been observed for the first time by laser inducedfluorescence (LIF) spectroscopy in the 480 - 430 nm region. HPBr and DPBr were generated in a pulsed discharge jet using a precursor mixture of PBr_{3} and H_{2} or D_{2} in argon. At low resolution the bands have the characteristic rotational contour of a perpendicular transition of an asymmetric top, and a band-dependent bromine isotope effect in an approximate 1:1 ratio of HP^{79}Br and HP^{81}Br. The analysis of the spectra, rotational constants, vibrational frequencies, molecular geometries and the results of ab initio calculations of the molecular parameters will be discussed.